how do you determine the length and width when you are given the difference of two similar objects
Answer by lynnlo(4176)

P=2L+2W
P=2(L+W)
P/2=L+W
IF THE PERIMETER IS GIVEN AND YOU KNOW ONE OF THE SIDES MEASURE LENGTH,THEN YOU CAN FIND THE MEASURE LENGTH OF THE SIDES,IF YOU KNOW THE MEASURE LENGTH OF L,THEN MEASURE LENGTH OF W IS P/2=W AND "VICE VERSA